# Exact Factorization of Unitary Transformations with Spin-Adapted Generators

### v1

@@ Appendix B Seniority 4, intermediate triplet
$A_{101}$ $=(\hat{n}(j_{\beta})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(j_{\beta})+\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{102}$ $=(\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{103}$ $=(\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})+2\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})$
$\quad-\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{104}$ $=(\hat{n}(b_{\beta})\hat{n}(a_{\beta})\hat{n}(j_{\beta})-\hat{n}(j_{\beta})\hat{n}(i_{\beta})-2\hat{n}(b_{\beta})\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\beta})+\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\beta})$
$\quad+\hat{n}(b_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(a_{\beta})+\hat{n}(b_{\beta})\hat{n}(a_{\beta})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{b_{\alpha}a_{\alpha}}$
$A_{105}$ $=(\hat{n}(j_{\beta})-\hat{n}(b_{\beta})\hat{n}(j_{\beta})-\hat{n}(a_{\beta})\hat{n}(j_{\beta})+\hat{n}(b_{\beta})\hat{n}(a_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{106}$ $=(\hat{n}(i_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{107}$ $=(\hat{n}(b_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\beta})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{108}$ $=(\hat{n}(b_{\beta})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-2\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})+\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})$
$\quad-\hat{n}(b_{\beta})\hat{n}(a_{\alpha})+\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{109}$ $=(\hat{n}(b_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(j_{\alpha})-\hat{n}(b_{\beta})\hat{n}(a_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{110}$ $=(\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})$
$\quad+\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(j_{\beta})+\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{111}$ $=(\hat{n}(a_{\beta})\hat{n}(j_{\beta})-\hat{n}(j_{\beta})\hat{n}(i_{\alpha})+\hat{n}(a_{\beta})\hat{n}(i_{\alpha})-\hat{n}(a_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{112}$ $=(\hat{n}(b_{\beta})\hat{n}(i_{\beta})+\hat{n}(a_{\alpha})\hat{n}(i_{\beta})-\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(a_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{113}$ $=(\hat{n}(i_{\beta})-\hat{n}(a_{\beta})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(i_{\beta})+\hat{n}(b_{\beta})\hat{n}(a_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{114}$ $=(\hat{n}(j_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{115}$ $=(\hat{n}(a_{\beta})-\hat{n}(a_{\beta})\hat{n}(j_{\beta})+\hat{n}(j_{\beta})\hat{n}(i_{\beta})-\hat{n}(a_{\beta})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{116}$ $=(\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(j_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{117}$ $=(\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(a_{\beta})\hat{n}(i_{\alpha})-\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{118}$ $=(\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})-\hat{n}(j_{\alpha})\hat{n}(i_{\beta})+\hat{n}(a_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{119}$ $=(\hat{n}(b_{\alpha})\hat{n}(j_{\beta})+\hat{n}(a_{\beta})\hat{n}(j_{\beta})-\hat{n}(j_{\beta})-\hat{n}(b_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{120}$ $=(\hat{n}(b_{\beta})\hat{n}(j_{\alpha})-\hat{n}(b_{\beta})-\hat{n}(j_{\alpha})\hat{n}(i_{\beta})+\hat{n}(b_{\beta})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{121}$ $=-(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(b_{\beta})+\hat{n}(j_{\beta})\hat{n}(a_{\alpha})$
@@
$A_{122}$ $=-(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\beta})\hat{n}(a_{\beta})+\hat{n}(i_{\beta})\hat{n}(b_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(a_{\beta})$
@@
$A_{123}$ $=(\hat{n}(b_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{124}$ $=(\hat{n}(b_{\beta})\hat{n}(a_{\alpha})+2\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})$
$\quad+\hat{n}(j_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{125}$ $=(\hat{n}(b_{\alpha})\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(j_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(i_{\beta})+2\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})$
$\quad+\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-\hat{n}(a_{\beta})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{126}$ $=(\hat{n}(a_{\beta})+\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-\hat{n}(a_{\beta})\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{127}$ $=(\hat{n}(b_{\beta})\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(i_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{128}$ $=(\hat{n}(b_{\beta})\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})\hat{n}(j_{\beta})-\hat{n}(a_{\alpha})\hat{n}(j_{\beta})+\hat{n}(j_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{129}$ $=(\hat{n}(b_{\beta})\hat{n}(j_{\beta})-2\hat{n}(b_{\beta})\hat{n}(a_{\beta})\hat{n}(j_{\beta})-2\hat{n}(b_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\beta})+4\hat{n}(b_{\beta})\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\beta})$
$\quad-2\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\beta})-2\hat{n}(b_{\beta})\hat{n}(a_{\beta})\hat{n}(i_{\beta})+\hat{n}(j_{\beta})\hat{n}(i_{\beta})+\hat{n}(a_{\beta})\hat{n}(i_{\beta})+\hat{n}(b_{\beta})\hat{n}(a_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{130}$ $=(4\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})$
$\quad-2\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})-2\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{131}$ $=(\hat{n}(b_{\alpha})\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})\hat{n}(j_{\alpha})+\hat{n}(j_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{132}$ $=(\hat{n}(b_{\alpha})\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})\hat{n}(i_{\beta})-\hat{n}(a_{\beta})\hat{n}(i_{\beta})+\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{133}$ $=(\hat{n}(j_{\beta})\hat{n}(i_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(j_{\beta})-\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{134}$ $=(\hat{n}(j_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(j_{\beta})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha})+\hat{n}(b_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{135}$ $=(2\hat{n}(a_{\alpha})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha})+2\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})$
$\quad-4\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})+2\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(b_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})$
$\quad-\hat{n}(a_{\alpha})\hat{n}(j_{\beta})-\hat{n}(b_{\beta})\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{136}$ $=(2\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})\hat{n}(j_{\alpha})+2\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})\hat{n}(i_{\beta})$
$\quad+2\hat{n}(a_{\beta})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-4\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-\hat{n}(b_{\alpha})\hat{n}(i_{\beta})$
$\quad-\hat{n}(b_{\alpha})\hat{n}(a_{\beta})+2\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{137}$ $=(2\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(j_{\alpha})\hat{n}(i_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-\hat{n}(a_{\alpha})\hat{n}(i_{\beta})$
$\quad-4\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(j_{\alpha})+2\hat{n}(b_{\beta})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})$
$\quad+2\hat{n}(b_{\beta})\hat{n}(a_{\alpha})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(a_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{138}$ $=(2\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})+2\hat{n}(b_{\alpha})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})\hat{n}(i_{\alpha})-4\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\alpha})$
$\quad+2\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(j_{\beta})-\hat{n}(b_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(i_{\alpha})$
$\quad-\hat{n}(a_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}.$
@@
Let us denote $\mathfrak{s}=[\mathcal{A}_{S4T},\mathcal{A}_{S4T}]$ a derived algebra that constitute the semi-simple part, $\dim\mathfrak{s}=132$. There is a 6 dimensional center, $Z(\mathcal{A}_{S4T})$. The solvable radical coincides with the center: $\mathfrak{r}=Z(\mathcal{A}_{S4T})$, and

### v2

@@ Appendix B Seniority 4, intermediate triplet
$A_{77}$ $=(\hat{n}(j_{\beta})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})\hat{n}(j_{\beta})+\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{78}$ $=(\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{79}$ $=(\hat{n}(b_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\beta})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{80}$ $=(\hat{n}(b_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(j_{\alpha})-\hat{n}(b_{\beta})\hat{n}(a_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{81}$ $=-(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(b_{\beta})+\hat{n}(j_{\beta})\hat{n}(a_{\alpha})$
@@
$A_{82}$ $=-(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\beta})\hat{n}(a_{\beta})+\hat{n}(i_{\beta})\hat{n}(b_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(a_{\beta})$
@@
$A_{83}$ $=(\hat{n}(b_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{84}$ $=(\hat{n}(b_{\beta})\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(i_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
Commuting groups
The groups of commuting generators are,
$s$ $=\{1,2,15,16,17,18,31,32,33,34,39,40,41,42,43,44,65,66,67,68,69,70,71,72,81,82,83,84\}$ (23)
$s$ $=\{3,4,9,10,13,14,19,20,23,24,27,28,45,46,47,48,51,52,55,56,57,58,61,62,73,74,79,80\}$ (24)
$s$ $=\{5,6,7,8,11,12,21,22,25,26,29,30,35,36,37,38,49,50,53,54,59,60,63,64,75,76,77,78\}$ (25)
@@
Let us denote $\mathfrak{s}=[\mathcal{A}_{S4T},\mathcal{A}_{S4T}]$ a derived algebra that constitute the semi-simple part, $\dim\mathfrak{s}=78$. There is a 6 dimensional center, $Z(\mathcal{A}_{S4T})$. The solvable radical coincides with the center: $\mathfrak{r}=Z(\mathcal{A}_{S4T})$, and